What are the IUPAC names of the molecules shown?


CH₃ -CH₂ - CH₂ - C = C-CH₃ С на сH3 СН2=СН-СН2-СН2-СН-СН3 3--3 CH₂
Answer #1

Answer) General Rules. - Select the longest chain containing double bond 09 CH₂=CH-CH2-CH2-CH-CH₂ CH₂ - CH₂ - CH₂-C=C-CH₂ che